# Diffusion as Reasoning: Enhancing Object Goal Navigation with LLM-Biased Diffusion Model

### v1

@@ Abstract
The Object Goal Navigation (ObjectNav) task requires the agent to navigate to a specified target in an unseen environment. Since the environment layout is unknown, the agent needs to perform semantic reasoning to infer the potential location of the target, based on its accumulated memory of the environment during the navigation process. Diffusion models have been shown to be able to learn the distribution relationships between features in RGB images, and thus generate new realistic images.In this work, we propose a new approach to solving the ObjectNav task, by training a diffusion model to learn the statistical distribution patterns of objects in semantic maps, and using the map of the explored regions during navigation as the condition to generate the map of the unknown regions, thereby realizing the semantic reasoning of the target object, i.e., diffusion as reasoning (DAR). Meanwhile, we propose the global target bias and local LLM bias methods, where the former can constrain the diffusion model to generate the target object more effectively, and the latter utilizes the common sense knowledge extracted from the LLM to improve the generalization of the reasoning process. Based on the generated map in the unknown region, the agent sets the predicted location of the target as the goal and moves towards it. Experiments on Gibson and MP3D show the effectiveness of our method.
Index Terms:
diffusion model; object goal navigation; semantic reasoning; common sense knowledge
I Introduction
@@
Embodied navigation, which requires an agent to use visual sensing to actively interact with its environment and perform navigation tasks, has seen rapid development driven by large-scale photorealistic domestic scene datasets [1], [2], [3] and advanced high-performance simulators [4], [5]. As a specific form of embodied navigation, Object Goal Navigation (ObjectNav) [6] is regarded as a key technology for enabling robotic intelligence. In ObjectNav task, the agent is placed in an unknown environment and is required to navigate to a user-specified object category (e.g. sofa) based on visual observations. Since the environment is unseen, when the target object is not visible, the agent needs to infer the potential locations of the target.
@@
The intelligence of an embodied navigation agent is primarily manifested in its semantic understanding and reasoning capabilities of the world (e.g., TVs are in the living room, ovens are in the kitchen, and chairs are near tables). Prior work has made some research progress in how to learn and incorporate the contextual relationships between objects into ObjectNav systems. Some methods extract relationship information among objects, regions, or rooms, and leverage this prior knowledge to enhance the visual representation of end-to-end navigation models [7], [8]. Other end-to-end approaches directly learn how to map the visual representation to actions by interacting with the environment [9], [10]. However, end-to-end methods implicitly embed semantic prior knowledge and simultaneously learn localization, mapping, and path planning in an implicit manner, which leads to the problems of high computational load and limited generalization in unseen environments.
Alternatively, modular methods tend to use a long-term goal policy module instead of the frontier exploration module, while also retaining the mapping and path planning modules in the traditional pipeline [11], [12], [13], [14]. However, the key issue remains how to represent and learn the contextual relations in the long-term goal policy module to perform semantic reasoning about the target location.
Some modular approaches use supervised learning to directly learn target relations, such as predicting the nearest potential frontier on a local map [12], estimating the distance to the target [15], or anticipating the target’s coordinates [16]. Other methods use the distance to the target as a reward to train modular-based reinforcement learning (RL) approaches [14]. However, both approaches fail to fully exploit the geometric and semantic cues in the map, as they implicitly and independently learn object relations. Therefore, learning the cooperative relationships among multiple objects in the scene has proven to be more reliable [13]. Furthermore, the ability of Large Language Models (LLMs) to extract and integrate knowledge for robotics has been widely studied [17]. The inherent common-sense knowledge in LLMs significantly improves the success rate of ObjectNav systems. However, zero-shot, training-free approaches [18] that rely solely on LLMs offer limited reliability when only sparse semantic information is available, as the reasoning capability of LLMs using their common-sense knowledge is diminished [19].
To alleviate the above problems, we propose an approach named DAR (Diffusion As Reasoning). This method combines the strengths of data-driven modular techniques with common-sense knowledge from LLMs about semantic and spatial relationships, enhancing the efficiency of ObjectNav systems.
The proposed DAR method uses Denoising Diffusion Probabilistic Models [20] (DDPM), conditioning on the agent’s memory of the environment to generate reliable semantic content in unknown areas. Following the self-supervised learning setup of DDPM, we replace the original 3-channel RGB data with semantic maps of the class channels, scaling up the denoising model accordingly. Our DAR is trained to predict pixel values in unexplored regions by leveraging both exploration memory and common sense knowledge. Exploration memory represents a local semantic map built from the agent’s visual observations of the environment. Unlike standard image datasets, ObjectNav has limited training scene quantity and diversity. To improve generalization, we also incorporate general knowledge from LLMs (e.g. GPT-4 [21]). Unlike existing method [22], the DAR approach does not require frequent LLM queries during navigation for valuable decisions. Instead, it relies on a knowledge graph, pre-extracted from LLMs, that represents object relationships.
In summary, our study has the following main contributions:
  • We use a DDPM to generate semantic maps, trained on public datasets [3], [23]. We propose a global target bias method to generate more specific types of objects, which controls the channel bias of the initial noise in the sampling process.
  • We pre-extract a knowledge graph from LLMs based on semantic and spatial relevance, and introduce a local LLM bias method to inject common-sense knowledge into the sampling process. This guides the DDPM to generate more specific objects in particular regions, according to the suggestions provided by the LLM.
  • We propose the DAR method as a modular-based solution to the ObjectNav problem. By using the trained denoising model, we integrate globally and locally biased initial noise and use an existing local semantic map as a condition to predict the object distribution in unexplored regions within a certain range. This enables semantic reasoning with a diffusion model, predicting potential locations of the target object.
II Related Work
II-A Object goal navigation
@@
Unlike end-to-end methods, modular approaches [12], [13] break down ObjectNav tasks into separate components: mapping [11], long-term goal reasoning [19], and path planning [32]. The mapping module provides a general representation that separates perception from reasoning and planning, making it more suitable for real-world applications.
@@
II-B Visual semantic reasoning
@@ II-C Diffusion models
III Methodology
III-A Task Definition
The ObjectNav task involves an agent navigating to an instance of a given object category (e.g., toilet) in an unknown environment. At the start of each episode, the agent is randomly placed. At each timestep $t$, it receives egocentric RGB-D observations, target object information, and its current pose ${\left(x,y,\theta\right)}$. The agent can perform one of four discrete actions: move forward, turn left, turn right, or stop. The agent autonomously executes a stop action when it determines the task is complete. An episode is considered successful if, within a given number of steps, the agent stops at a position where the distance to the target is below a threshold (e.g., 1m) and the target is visible in its egocentric view.
Our method builds on the modular ObjectNav architecture, which typically constructs an accumulating semantic map during navigation The semantic map (${m_{t}\in\mathbb{R}^{\left(N_{o}+N_{s}\right)\times H\times W}}$) uses multiple channels to represent different classes: ${N_{o}}$ for occupancy classes (occupied and free), and ${N_{s}}$ for object classes, with $H$ and $W$ as the map dimensions. This map accumulates observed object layouts of the environment from timesteps 0 to $t$.
@@ III-A Task Definition
III-B Self-supervised Diffusion Learning
In the field of RGB image generation, trained diffusion models can learn the latent statistical distribution of image features. We have observed that the spatial distribution of objects in indoor scenes also follows certain statistical patterns. Since diffusion models can generate high-quality RGB images, they should, with sufficient training, also be able to generate high-quality maps of room object distributions. Building on this idea, we used a room semantic map dataset to train an unconditional diffusion model.
@@ III-B Self-supervised Diffusion Learning
If the variance schedule ${\left(\beta_{t}\right)_{t=1}^{T}}$ makes ${\hat{\beta}_{T}\rightarrow 0}$, then ${x_{T}}$ approaches the standard Gaussian distribution, namely ${x_{T}\rightarrow\mathcal{N}\left(0,\mathbf{I}\right)}$. The training process is represented by reversing the forward diffusion process. By training a neural network model to predict the noise corresponding to each time step, the noise can then be removed one by one, gradually recovering the original data ${x_{0}}$. This process is described as Eq. 2, where $\theta$ represents the parameters of the neural network model, and the network takes noisy data ${x_{t}}$ and time step $t$ as input to predict the mean ${\mu_{\theta}\left(x_{t},t\right)}$ and covariance ${\Sigma_{\theta}\left(x_{t},t\right)}$.
${p_{\theta}\left(x_{t-1}\middle|x_{t}\right)=\mathcal{N}\left(x_{t-1};\mu_{\theta}\left(x_{t},t\right),\Sigma_{\theta}\left(x_{t},t\right)\right)}$ (2)
Ho et al. [20] proposed a simplified method, where the covariance ${\Sigma_{\theta}\left(x_{t},t\right)}$ is fixed to a constant value and the mean ${\mu_{\theta}\left(x_{t},t\right)}$ is rewritten as a function of noise, as follows:
${\mu_{\theta}=\frac{1}{\sqrt{\alpha_{t}}}\cdot\left(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\hat{\beta}_{t}}}\cdot z_{\theta}\left(x_{t},t\right)\right)}$ (3)
@@
${\mathcal{L}_{vlb}=\mathbb{E}_{t\sim\left[1,T\right]}\mathbb{E}_{x_{0}\sim p\left(x_{0}\right)}\mathbb{E}_{z_{t}\sim\mathcal{N}\left(0,\mathbf{I}\right)}\left\|z_{t}-z_{\theta}\left(x_{t},t\right)\right\|^{2}}$ (4)
Figure 1: Diffusion learning process.
@@
Where $\mathbb{E}$ is the expected value. As a result, the neural network model no longer needs to predict both the mean and covariance simultaneously, but only needs to estimate the noise ${z_{\theta}\left(x_{t},t\right)}$ corresponding to ${x_{t}}$.
@@
As shown in Figure 1, the input ${x_{0}}$ contains 18 binary semantic category channels (using the GIBSON dataset as an example). Additionally, we rescale the room semantic maps from the dataset to a resolution of ${256\times 256}$. Following the training process of diffusion models for RGB images, the model architecture is consistent with guided-diffusion [37], with increased depth and channel count of the U-Net model. The trained diffusion model can generate a regularized RGB image from noise. Similarly, by using 18-channel Gaussian noise as input, the model progressively removes the noise and ultimately generates a semantic map. Note that the output of the denoising model does not produce a standard multi-channel binary-valued semantic map. Therefore, we use the argmax operation to post-process the results and obtain the output shown in Figure 3.
Figure 2: The effect of global and local biases.
Figure 3: The effect of global target bias.
Next, we raise another question: How can we make the generated semantic maps from the diffusion model contain more ’specified’ objects? This problem is similar to the field of conditional image generation, but it is not feasible to train a conditional denoising model using classification ideas on multi-channel semantic maps, as the meaning of the map lies in the distribution of the objects it contains, and considering the overall category of the map is meaningless. In this paper, we propose a simple and effective way to address this question.
For the target channel, the noise is replaced with a biased Gaussian noise distribution with mean ${\bm{\mu}_{b}}$ and covariance ${\mathbf{\Sigma}_{b}}$, while the other channels remain unchanged. We refer to this process as ’global target bias’. Compared to using standard Gaussian noise across all channels, using the target-biased noise as the initial input results in the denoising model generating outputs containing more objects corresponding to the target channel.
As shown in Figure 3, applying bias to the TV channel during the sampling process results in more intermediate noise representing TV, visible as green pixels in sections ${x_{T-1}^{b}}$ and ${x_{40}^{b}}$. This eventually leads to the generation of a map containing more TV instances, which are marked with red circles in Figure 3. In contrast, the map generated without bias during sampling does not contain any TV at all.
The two sampling processes shown in Figure 3 used the same torch random seed to demonstrate the effectiveness of the global target bias. Additionally, even with the bias towards TV channel, the model did not randomly generate a chaotic TV grid, but rather generated the TV near book, chair, and couch, which aligns with the actual patterns. This suggests that after the global target bias, the model was able to generate more of the target objects according to the learned distribution of room objects. More examples of global target bias generation can be seen in Figure 2. All the maps were generated by the diffusion model.
III-C Semantic Reasoning with Local Map Condition

### v2

@@ Abstract
The Object Navigation (ObjectNav) task aims to guide an agent to locate target objects in unseen environments using partial observations. Prior approaches have employed location prediction paradigms to achieve long-term goal reasoning, yet these methods often struggle to effectively integrate contextual relation reasoning. Alternatively, map completion-based paradigms predict long-term goals by generating semantic maps of unexplored areas. However, existing methods in this category fail to fully leverage known environmental information, resulting in suboptimal map quality that requires further improvement. In this work, we propose a novel approach to enhancing the ObjectNav task, by training a diffusion model to learn the statistical distribution patterns of objects in semantic maps, and using the map of the explored regions during navigation as the condition to generate the map of the unknown regions, thereby realizing the long-term goal reasoning of the target object, i.e., diffusion as reasoning (DAR). Meanwhile, we propose the Room Guidance method, which leverages commonsense knowledge derived from large language models (LLMs) to guide the diffusion model in generating room-aware object distributions. Based on the generated map in the unknown region, the agent sets the predicted location of the target as the goal and moves towards it. Experiments on Gibson and MP3D show the effectiveness of our method.
keywords:
[…]
1 Introduction
@@
Embodied navigation, which requires an agent to use visual sensing to actively interact with its environment and perform navigation tasks, has seen rapid development driven by large-scale photorealistic domestic scene datasets [4], [33], [46] and advanced high-performance simulators [9], [37]. As a specific form of embodied navigation, Object Navigation (ObjectNav) [22] is regarded as a key technology for enabling robotic intelligence. In ObjectNav task, the agent is placed in an unknown environment and is required to navigate to a user-specified object category (e.g., sofa) based on visual observations. Since the environment is unseen, when the target object is not visible, the agent needs to infer the potential locations of the target.
@@
Embodied navigation agents show intelligence through semantic understanding and reasoning. Previous work explored learning object contextual relationships for ObjectNav systems. Some methods extract object/region/room relationships to enhance visual representation [13], [12]. while other end-to-end approaches directly map vision to actions [7], [30]. However, end-to-end methods implicitly embed semantic prior knowledge and simultaneously learn localization, mapping, and path planning in an implicit manner, which leads to the problems of high computational load and limited generalization in unseen environments.
Alternatively, modular approaches typically address the ObjectNav problem by incorporating three modules: Mapping, Long-term Goal Policy, and Path Planning. Among these, the Long-term Goal Policy module plays a pivotal role. When the target object is absent from the currently constructed local map, a semantic reasoning algorithm is employed to infer an appropriate location as the long-term goal, which subsequently serves as input to the Path Planning module. Notably, the accuracy of the Long-term Goal Policy critically determines both the success rate and efficiency of the ObjectNav system.
The existing approaches for long-term goal inference can be broadly categorized into two classes: location prediction methods and map completion methods.
Location prediction methods refer to algorithms that directly predict the position of the long-term goal based on the semantic map constructed from explored regions and the target object category. Some methods perform location prediction along the frontiers of the local map, selecting the point closest to the target object as the long-term goal [32], or incorporating common-sense knowledge to enhance frontier-based predictions [43]. Other methods extend beyond frontier regions, instead predicting the target object’s location globally and designating it as the long-term goal [5] [52]. Additionally, Xinyao Yu et al. proposed a trajectory prediction approach [51], which leverages the agent’s historical observations and the target object as conditions to predict a sequence of long-term goals. This method, in essence, also falls under the category of location prediction.
However, location prediction methods suffer from two key limitations: (1) their reliance on pre-constructed datasets, and (2) their individual learning of target-object relations, which may vary significantly across different room layouts. These limitations are effectively addressed by map completion methods, which adopt a self-supervised learning paradigm to jointly model inter-object relationships (referred to as contextual object relations). Recent advances in map completion have employed encoder-decoder architectures to model these contextual object relations.
Several approaches employ CNN-based models (e.g., U-Net [36]) to predict unobserved regions based on single-timestamp top-down maps [25] [18]. Furthermore, SGM [54] adopts a masked modeling strategy [20] to train a transformer-based model (ViT [11]), which integrates visual observations with LLM knowledge (e.g., GPT-4 [1], ChatGLM [15]). This framework operates by first patchifying the input map and then predicting missing patch values conditioned on visible patches.
The map completion method predicts long-term goals by generating semantic maps of unknown regions. Undoubtedly, the higher the reliability of the generated maps, the greater the improvement in the performance of the ObjectNav system. However, the quality of the unknown-area maps generated by the aforementioned methods remains suboptimal. On one hand, these methods focus solely on objects themselves, while the absolute spatial positions of objects may vary across different environments. On the other hand, existing methods fail to fully utilize the information from observed regions.
We observe that learning contextual object relations equates to learning the statistical distribution patterns of object placements within indoor rooms. Generating the object distribution maps in unknown regions equates to sampling from the distributions that best match those in the observed areas.
Inspired by the remarkable success of diffusion models in representing complex distributions [21], we propose replacing conventional CNN-based and Transformer-based architectures in self-supervised map generation with a diffusion model framework. Unlike existing approaches that focus solely on object-level features, our method incorporates room category information as conditional guidance. While object locations may exhibit ambiguity across different room layouts, specific room categories (e.g., bedrooms) consistently contain characteristic object classes (e.g., beds, pillows). By leveraging prior knowledge of a region’s room category, our model can more effectively constrain the generation process to avoid implausible object distributions, thereby significantly improving the quality of generated maps for unknown regions.
In this paper, we present a modular solution for the ObjectNav problem, introducing a Diffusion-based Approach for long-term goal Reasoning (DAR). Our framework conditions on the agent’s environmental memory while leveraging room category distributions as guidance to generate reliable semantic content in unknown areas. The DAR module selects potential target object locations within these generated regions as long-term goals. For training DAR, we utilize standard ObjectNav datasets (Gibson, Matterport3D, etc.) to obtain ground-truth semantic room maps. These maps undergo random rotations, mirroring, and translations to construct the training set for our diffusion model. Following the self-supervised learning paradigm of DDPM, our DAR model learns to generate plausible, novel room semantic maps - distinct from training samples - through iterative denoising from pure Gaussian noise.
However, unlike image datasets, training scenes for ObjectNav suffer from notable limitations in both quantity and diversity. To enhance generalization capability, previous approaches have incorporated general knowledge from LLMs. Our DAR module adopts this paradigm through object-room relationship modeling. The denoising model in DAR takes both the denoised result xt from the previous time step and a room guidance map as input. This approach enables the diffusion model to generate specific object clusters according to room types. Since diffusion models have demonstrated the capability to generate pixel values in unknown regions based on partially known areas without requiring retraining, we leverage this mechanism by using the semantic map of known regions from actual navigation as a condition. The pre-trained diffusion model is employed to generate semantic maps for unknown regions, with room type guidance also applied. We first determine the room type of frontier regions using common-sense knowledge extracted from LLMs, based on the observed object distribution in the established semantic map. For instance, if objects near frontier point $p$ include oven, sink, and refrigerator ordered by proximity, the LLM would identify $p$ as located in a kitchen. This room type then guides the diffusion model to generate object distributions around $p$ that are consistent with a kitchen (e.g., cup, bottle) rather than incompatible objects (e.g., bed, tv).
We evaluate our DAR model in simulation environments, including photorealistic 3D platforms: AI2-Thor, Robo-Thor, Gibson, and MP3D. Experimental results demonstrate that our DAR model significantly outperforms previous map completion methods, achieving substantial improvements in map quality, object diversity, and room structure/contour accuracy. The precise semantic understanding enables DAR to produce more accurate semantic maps, leading to better long-term goal prediction and more efficient object navigation.
In summary, our study makes the following key contributions:
  • We propose a modular-based ObjectNav approach, DAR, that learns statistical distributions of rooms and objects, enabling accurate generation of unknown-region object layouts based on explored-area maps.
  • Leveraging room-object correlations extracted from LLMs, we infer potential room types at frontier regions of partial semantic maps.
  • We design a denoising model augmented with a room-type guidance submodule, ensuring generated object distributions adhere to inferred spatial semantics.
  • Extensive evaluations on Gibson and Matterport3D (MP3D) demonstrate that DAR achieves state-of-the-art Success Rate (SR). Notably, it improves SPL (efficiency metric) by 10.6% on average over prior works across both datasets.
2 Related Work
2.1 Object navigation
@@
Unlike end-to-end methods, modular approaches [32], [54] break down ObjectNav tasks into separate components: mapping [5], long-term-goal reasoning [43], and path planning [38]. The mapping module provides a general representation that separates perception from reasoning and planning, making it more suitable for real-world applications.
@@
2.2 Visual semantic reasoning
@@ 2.3 Diffusion models
3 Methodology
3.1 Task Definition
The ObjectNav task involves an agent navigating to an instance of a given object category in an unknown environment. At the start of each episode, the agent is randomly placed. At each timestep $t$, it receives egocentric RGB-D observations, target object information, and its current pose ${\left(x,y,\theta\right)}$. The agent can perform one of four discrete actions: move forward, turn left, turn right, or stop. The agent autonomously executes a stop action when it determines the task is complete. An episode is considered successful if, within a given number of steps, the agent stops at a position where the distance to the target is below a threshold (e.g., 1m) and the target is visible in its egocentric view.
Our method builds on the modular ObjectNav architecture, which typically constructs an accumulating semantic map (${m_{t}\in\mathbb{R}^{\left(N_{o}+N_{s}\right)\times H\times W}}$) during navigation. The semantic map uses multiple channels to represent different classes: ${N_{o}}$ for occupancy classes (occupied and free), and ${N_{s}}$ for object classes, with $H$ and $W$ as the map dimensions. This map accumulates observed object layouts of the environment from timesteps 0 to $t$.
@@ 3.1 Task Definition
3.2 Self-supervised Diffusion Learning
3.2.1 Data Preparation
The DAR model is trained in a self-supervised manner. We extract semantic maps for each floor of every scene from the standard ObjectNav dataset. To obtain ground-truth room segmentation maps, we note that both the Gibson [46] and Matterport3D [4] datasets provide annotated object and room/area categories for each house (e.g., the 3D scene graph [2] with Gibson annotations). Following the same processing method as LROGNav [43], we compute 3D bounding boxes for each room using face IDs and key house parameters (e.g., floors and floor heights), then project the house point cloud into 2D floor maps to derive top-down room segmentation, as shown in Fig. 1(b).
Figure 1: (a) Object semantic map, color encoded; (b) Room semantic segmentation map; (c) Room guidance with points; (d) Room guidance with exponentially decreasing distribution.
So far, taking Gibson as an example, we have obtained an $H\times W\times 18$ object map $M_{o}$ (with 18 object categories, Fig. 1(a)) and an $H\times W\times 19$ room segmentation map $M_{r}$ (with 19 room categories, Fig. 1(b)). The DAR model aims to leverage the room segmentation map as guidance in the reverse process, where $M_{r}$ should provide prior information about the distribution of objects within certain regions. However, the current $M_{r}$ contains clear boundaries between different rooms, which, while reasonable in the dataset, may not hold in real-world navigation—agents can hardly infer exact room boundaries from limited observations. Thus, we want $M_{r}$ to offer object distribution priors without imposing structural constraints on the generated object map.
To address this, we modify $M_{r}$ as follows: For each room category channel, we first cluster regions of that category, then compute the geometric center ($c_{x}$, $c_{y}$) of each cluster. Finally, we design a distribution map where pixel values decay exponentially from the center outward, as defined in Eq. 1.
$M_{r}^{i}\left(x,y\right)=\exp\left(-\frac{\sqrt{\left(x-c_{x}\right)^{2}+\left(y-c_{y}\right)^{2}}}{\sigma^{2}}\right)$ (1)
Here, $M_{r}^{i}$ denotes the i-th channel of the room segmentation map, where $(x,y)$ represents pixel coordinates. For each room cluster, we compute its minimum bounding box and set $\sigma$ (the decay coefficient) as one-fourth of the average side length of the bounding box, and get room guidance as shown in Fig. 1(d). In Section 4.2, we evaluate how different approaches for determining the $\sigma$ value affect the quality of the generated maps in unexplored regions.
3.2.2 Denoising Model Design
We employ a U-Net-based architecture as our backbone model and investigate three distinct denoising network configurations for the diffusion process, as shown in Fig. 5.
  • Concatenation: The most straightforward approach directly concatenates the input $m_{r}$ and $x_{t}$ before feeding them into the U-Net denoising network. While simple to implement, this method significantly increases the channel dimension (to objects_class_num + rooms_class_num), potentially affecting computational efficiency and model performance.
  • Cross-Attention: We adopt the same architecture as the U-Net encoder blocks to construct the Room Guidance encoder blocks. For notational convenience, the output of $x_{t}$ processed through the U-Net encoder blocks is denoted as $x_{en}$, while the output of $m_{r}$ processed through the Room Guidance encoder blocks is denoted as $m_{en}$. $x_{en}$ and $m_{en}$ undergo linear transformations via three separate $1\times 1$ convolutional layers to generate the query $q$, key $k$, and value $v$, respectively. The output is computed using Eq. 2, and the cross-attention mechanism is illustrated in Fig. 2.
    $\mathrm{CrossAttention}\left(q,k,v\right)=\mathrm{softmax}\left({\frac{qk^{T}}{\sqrt{d_{k}}}}\right)v$ (2)
  • Multi-scale Hierarchical Fusion and Semantic Attention: Inspired by RefineCatDiff [26], we construct the Room Guidance Encoder using a Conv-LayerNorm-ReLU architecture and extract multi-scale features from each layer, denoted as $r_{e}$, ensuring that its channel dimension matches that of the corresponding U-Net layer feature $x_{e}$. Subsequently, feature fusion is performed using Eq. 3, where LN denotes LayerNorm, RF denotes Room Feature Fusion as in Fig. 5.
    $\begin{cases}f=\mathrm{Concat}\left(x_{e},r_{e}\right)\\ f^{\prime}=\mathrm{ReLU}\left[\mathrm{LN}\left(\mathrm{Conv}\left(f\right)\right)\right]\\ \mathrm{RF}\left(x_{e},r_{e}\right)=\mathrm{LN}\left(r_{e}\oplus f^{\prime}\right)\\ \end{cases}$ (3)
    To effectively inject high-level semantic features from the room guidance into the network, we design a Semantic Attention (SA) module that integrates both channel attention and spatial attention mechanisms. The SA module consists of parallel spatial and channel attention branches that jointly process $r_{e}$ and $x_{e}$, generating channel attention weights ($\alpha_{ch}$) and spatial attention weights ($\alpha_{sp}$), as shown in Fig. 3. Each branch performs element-wise multiplication between the input features and their respective attention weights. The refined features are then combined through element-wise addition, and the enhanced output is produced via a convolutional layer for further processing.
Figure 2: Illustration of the cross-attention mechanism.
Figure 3: Structure of Semantic Attention module.
3.2.3 Diffusion Learning
@@
If the variance schedule ${\left(\beta_{t}\right)_{t=1}^{T}}$ makes ${\hat{\beta}_{T}\rightarrow 0}$, then ${x_{T}}$ approaches the standard Gaussian distribution, namely ${x_{T}\rightarrow\mathcal{N}\left(0,\mathbf{I}\right)}$. The training process is represented by reversing the forward diffusion process. By training a neural network model to predict the noise corresponding to each time step, the noise can then be removed one by one, gradually recovering the original data ${x_{0}}$. This process is described as Eq. 5, where $\theta$ represents the parameters of the neural network model, and the network takes noisy data ${x_{t}}$ and time step $t$ as input to predict the mean ${\mu_{\theta}\left(x_{t},t,r\right)}$ and covariance ${\Sigma_{\theta}\left(x_{t},t,r\right)}$. Since our denoising process incorporates room guidance, the denoising model takes an additional input, denoted as $r$.
${p_{\theta}\left(x_{t-1}\middle|x_{t}\right)=\mathcal{N}\left(x_{t-1};\mu_{\theta}\left(x_{t},t,r\right),\Sigma_{\theta}\left(x_{t},t,r\right)\right)}$ (5)
Ho et al. [21] proposed a simplified method, where the covariance ${\Sigma_{\theta}\left(x_{t},t,r\right)}$ is fixed to a constant value and the mean ${\mu_{\theta}\left(x_{t},t,r\right)}$ is rewritten as a function of noise, as follows:
${\mu_{\theta}=\frac{1}{\sqrt{\alpha_{t}}}\cdot\left(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\hat{\beta}_{t}}}\cdot z_{\theta}\left(x_{t},t,r\right)\right)}$ (6)
@@
${\mathcal{L}_{vlb}=\mathbb{E}_{t\sim\left[1,T\right]}\mathbb{E}_{x_{0}\sim p\left(x_{0}\right)}\mathbb{E}_{z_{t}\sim\mathcal{N}\left(0,\mathbf{I}\right)}\left\|z_{t}-z_{\theta}\left(x_{t},t,r\right)\right\|^{2}}$ (7)
Figure 4: Diffusion model learning process based on semantic maps, incorporating room guidance during denoising.
@@
Where $\mathbb{E}$ is the expected value. As a result, the neural network model no longer needs to predict both the mean and covariance simultaneously, but only needs to estimate the noise ${z_{\theta}\left(x_{t},t,r\right)}$ corresponding to ${x_{t}}$.
@@
As shown in Fig. 4, the input ${x_{0}}$ contains 18 binary channels (using the Gibson dataset as an example). Following the standard training procedure of diffusion models for RGB images, the trained model can generate a denoised RGB image from noise. Similarly, when provided with 18-channel Gaussian noise as input, the model iteratively denoises the signal and eventually produces a semantic map. Note that the model’s output is not a standard multi-channel binary semantic map by default. Thus, we apply an argmax operation as post-processing to discretize the predictions into valid semantic labels.
Figure 5: Denoising model architecture. Left: We train a conditional denoising network with three inputs, outputting noise and variance. Right: Detailed model structure. We design and study three variants, incorporating room feature fusion, cross-attention mechanisms, and simple concatenation, respectively.
3.3 Unobserved regions generation Guided by LLM-Based Room Priors
We constructed a dataset in Section 4.2.1 to validate the map completion model. In this section, an example from this dataset is used for illustration. As shown in Fig. 6, an object semantic map of the observed area was built during a navigation instance, with our goal being to generate the map content for unobserved regions. We first describe how to obtain LLM-based room priors for frontiers and finally explain how to generate the unknown area map using the pre-trained model from Section 3.2.
Figure 6: Room guidance computation pipeline.
3.3.1 LLM-based Object-Room Priors
To improve ObjectNav task efficiency, a key aspect is equipping the robot with semantic reasoning capability. Previous position prediction methods employed LLM-based object-to-object [50], room-to-room [17], and object-to-room [43] relationships to enhance navigation accuracy and generalization in unseen scenarios. However, these approaches applied commonsense knowledge to train the long-term goal prediction module, learning target relations individually. In contrast, we utilize LLM-based knowledge to guide map generation in unknown areas, employing object-room priors to direct a diffusion model in jointly learning multi-object relations—a method we demonstrate to be more reliable and efficient.
Similar to LROGNav [43] and LFG [39], we employ positive and negative prompts and utilize a chain-of-thought [44] approach to query the large language model (LLM) for determining the likelihood of each object belonging to specific room types. This yields an $N\times M$ matrix $V$, where $N$ and $M$ represent the number of object and room categories, respectively. Here, $V_{j}\left(o\right)$ denotes the probability of object $o$ belonging to room type $j$. Next, our objective is to estimate the room categories associated with frontiers in the local map. Following PONI [32], we define a frontier as a non-obstacle region at the boundary between known and unknown areas.
As shown in the Fig. 6, for frontier $F$, we select $n$ objects present in the map $\left[o_{1},o_{2},...,o_{n}\right]$ ($n=4$ in this case) and compute their corresponding weights $\left[w_{1},...,w_{n}\right]$ based on their distances to objects in the local map, where $\sum\nolimits_{1}^{n}{w_{i}}=1$. We then compute the probability of frontier $F$ belonging to room category $j$ using Eq. 8.
$R_{j}\left(F\right)=\sum\nolimits_{i=1}^{n}{V_{j}\left(o_{i}\right)\cdot w_{i}}$ (8)
Specifically, for each of the $n$ adjacent objects near $F$, we calculate its probability of belonging to room $j$ and multiply it by the corresponding weight. he final room assignment is determined by selecting the category with the highest probability, as formalized in Eq. 9.
$R(F)=\underset{j\in 1,...,M}{\mathrm{argmax}}R_{j}(F)$ (9)
Next, we employ Eq. 1 to design a distribution map within each channel, where pixel values decay exponentially from the center outward, as illustrated in Fig. 6 .
3.3.2 Map-Completion via Local Map-Conditioned Generation